OpenStudy (anonymous):

Identify the type of conic section consisting of the set of all points in the plane for which the absolute value of the difference of the distances from the points (0, 9) and (0, -9) is 3.

OpenStudy (anonymous):

|√(x² + (y - 9)²) - √((x² + (y + 9)²)| = 3

OpenStudy (anonymous):

√(x² + (y - 9)²) - √(x² + (y + 9)²) = ±3 x² + (y - 9)² - 2√((x² + (y - 9)²)(x² + (y + 9)²)) + (x² + (y + 9)²) = 9 x² + (y - 9)² + (x² + (y + 9)²) - 9 = 2√((x² + (y - 9)²)(x² + (y + 9)²)) (x² + (y - 9)² + (x² + (y + 9)²) - 9)² = 4(x² + (y - 9)²)(x² + (y + 9)²)